OpenStudy (anonymous):

the hypotenuse of a right triangle is 25 cm, and thr shorter leg is 15 cm. Find the length of the other leg.

OpenStudy (anonymous):

you need to use the Pythagorean Theorem A^2+ B^2= C^2 the hypotenuse is always across from the right angle, and is always C so.... A^2+ 15^2= 25^2 A^2+225=625 -225 -225 _______________ A^2= 400 then you square root both sides A= 20cm/short leg :)

OpenStudy (anonymous):

long leg not short leg srry :)

OpenStudy (anonymous):

thanks! can u help me more here?

OpenStudy (anonymous):

sure what do you need?

OpenStudy (anonymous):

which of the following equation is of a parabola with a vertex at (0,-5)? y=(x- 5) 2^, y= ( x+5)2^, y+ x2^-5, or y+ x 2^+5?

OpenStudy (anonymous):

what of the circle that has a diameter endpoints are (2,7) and -6-1)

OpenStudy (anonymous):

find the distance between the points (-4, -5) and (3,-1.

OpenStudy (anonymous):

the first Q is y= (x+5)^2

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

for the second one im not quite sure what your asking...are you asking what is the center called? or what...sorry

OpenStudy (anonymous):

what of the circle that has a diameter endpoints are (2,7) and -6-1). that one? yes.

OpenStudy (anonymous):

yeah that one

OpenStudy (anonymous):

k, yes that's what i'm asking

OpenStudy (anonymous):

different question: what is the center of the circle that has a diameter whose endpoints are (2,7) and (-6, -1)? ans: (4,3), (-2, 4), or (-2,3)?

OpenStudy (anonymous):

okee for that one you set it up like this (X1+X2), (Y1+Y2) _______ _______ 2 2 so (2-6, 7-1) ____ ____ 2 2 (-4, 6) ___ ___ 2 2 so therefore the answer is (-2, 3)

OpenStudy (anonymous):

the distance one you use the distance formula D=√(X2-X1)^2 + (Y2-Y1)^2 D=√(3+4)+(-1+5)= D=√(8)^2+ (4)^2= D=V64+16= D=8.9

OpenStudy (anonymous):

What question is that answer for?

OpenStudy (anonymous):

find the distance between the points (-4,-5), (3,-1)